OpenStudy (anonymous):

can anyone help me with 9th grade math?

OpenStudy (anonymous):

\[5/3 + 5 -5 = 8 - 5\]

OpenStudy (anonymous):

Isn't true what you have written there is 5/3 = 3

OpenStudy (anonymous):

no its write the way i typed it

OpenStudy (anonymous):

there are no variables in it

OpenStudy (anonymous):

the 5/3 is sapose to be n/3

OpenStudy (anonymous):

ah ok \[n/3 + 5 -5 =8 -5 \] \[n/3 +(5-5) =(8-5)\] \[n/3+0 = 3\] \[n/3 = 3\] \[n=3\times3=9\]

OpenStudy (anonymous):

thanks can u help me on some more cuz i dnt understand this stuff

OpenStudy (anonymous):

\[n/7 + 6 = 8 \]

OpenStudy (anonymous):

\[n/7 = 8-6 =2\] \[n=14\]